ce aYe reedyce sheu a second humb er)
Answer #1

Answer :

The number that comes of first should be anything, then we want it to be on next two throws , provided the other two throws are same. Or iit to be on next one throw provided the other three throws are different.

Probability : = (1/6)^2(1/5)^2 *6 + (1/6) *(1/5) ^3 *4 = 1/150 + 4/750 = 0.012